OpenStudy (openstudyrocks5*):

Solve The Equation. x/4 (meaning a fraction) + 1 = -5. Please Help

OpenStudy (anonymous):

First step is to subtract 1 to both sides. -5 - 1 = ?

OpenStudy (anonymous):

\[\frac{ x }{ 4 }+1=-5\]

OpenStudy (anonymous):

First step subtract 1 on both sides second step,multiply 4 on both sides

OpenStudy (anonymous):

-5-1=?

OpenStudy (anonymous):

Yep^^

OpenStudy (anonymous):

\[\frac{ x }{ 4 }+1-1=-5-1\]

OpenStudy (anonymous):

\[\frac{ x }{ 4 }=-6\]

OpenStudy (anonymous):

\[(\frac{ x }{ 4 })4=-6\times4\]

OpenStudy (anonymous):

\[x=?\]solve it @OpenStudyRocks5*

OpenStudy (openstudyrocks5*):

x = -24? @MARC_

OpenStudy (anonymous):

yep.^_^

OpenStudy (openstudyrocks5*):

Thank you. I may have more questions.

OpenStudy (anonymous):

yw!